OpenStudy (anonymous):

A customer from Cavallaro's Fruit Stand picks a sample of 5 oranges at random from a crate containing 65 oranges, of which 6 are rotten. What is the probability that the sample contains 1 or more rotten oranges? (Round your answer to three decimal places.)

OpenStudy (anonymous):

"one or more" makes it easy compute the probability that you get zero rotten ones subtract that number from 1

OpenStudy (anonymous):

you know how to do that?

OpenStudy (anonymous):

if not let me know and i can walk you through it just asking

OpenStudy (anonymous):

please walk me through it

OpenStudy (anonymous):

we are going to compute the probability that NONE are rotten first one is not probability is \(\frac{59}{65}\) since there are 59 not rotten ones and 65 all together

OpenStudy (anonymous):

second one is not rotten given the first one is not that is \(\frac{58}{64}\) because we know the first one selected was not rotten leaving 58 ok ones and 64 total because we picked one already

OpenStudy (anonymous):

and so one for all 5 not being rotten in other words \[\frac{59}{65}\times\frac{58}{64}\times \frac{57}{63}\times \frac{56}{62}\times \frac{55}{61}\]

OpenStudy (anonymous):

that is the probability that none of the five chosen are rotten the probability that one or more is is \(1-\frac{59}{65}\times\frac{58}{64}\times \frac{57}{63}\times \frac{56}{62}\times \frac{55}{61}\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

that is all

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw